# ProDAG: Projected Variational Inference for Directed Acyclic Graphs

### v5

@@ Appendix B Projection algorithms
We take $\mu^{(1)}=1$, $\mu^{(t+1)}=\mu^{(t)}/2$, $T=10$, and $W^{(0)}=0\in\mathbb{W}$ in the experiments.
Algorithm 3 provides the routine for the projection of the matrix $\hat{W}$ onto the $\ell_{1}$ ball of size $\lambda$.
@@
  Sort $v=\operatorname{vec}(\hat{W})$ as $|\mathsf{v}_{1}|\geq|\mathsf{v}_{2}|\geq\cdots\geq|\mathsf{v}_{p\times p}|$
@@ Appendix E Implementation
  • DiBS and DiBS+: does not include a sparsity parameter;
@@
The experiments are run on a Linux workstation with an AMD Ryzen Threadripper PRO 5995WX CPU, 256GB RAM, and 2 × NVIDIA GeForce RTX 4090 GPUs. Each method is allocated either a single GPU or a single CPU core, with the experiments run in parallel across GPUs/cores.
@@ Appendix F Simulation design
The synthetic datasets generated from linear DAGs are constructed as follows. We first randomly sample an Erdős–Rényi or scale-free graph with $p$ nodes and $s$ edges. The edges of this undirected graph are then oriented according to a randomly generated topological ordering. Each directed edge in the graph is then weighted by sampling weights uniformly on $[-0.7,-0.3]\cup[0.3,0.7]$, yielding a weighted adjacency matrix $W$. We then draw iid noise $\varepsilon_{1},\dots,\varepsilon_{n}$ and generate the variables as
@@
For synthetic datasets generated from nonlinear DAGs, we likewise sample an Erdős–Rényi or scale-free graph and orient it according to a random topological ordering. We then generate a feedforward neural network $f_{k}$ with a single hidden layer of 10 neurons for $k=1,\dots,p$. The weights of $f_{k}$ are sampled uniformly from $[-0.7,-0.3]\cup[0.3,0.7]$, with hidden layer weights of node $k$’s non-parents set to zero. The activation is a ReLU. The noise $\varepsilon_{1},\dots,\varepsilon_{n}$ is drawn iid and the variables are taken as
@@ Appendix G Additional experiments
We investigate the robustness of ProDAG to several alternative simulation designs. Figures 6 and 7 report results on datasets simulated from dense graphs containing two and three times as many edges as used in the main experiments. Figure 8 and 9 report the results from non-Gaussian noise distributions, specifically Gumbel and exponential distributions. Finally, Figure 10 and 11 report results on scale-free graphs instead of Erdős–Rényi graphs, where the degree exponent is set to 2 and 3. Under all three simulation designs, the relative performance of ProDAG is roughly similar to that in the main experiments, indicating that ProDAG is reliable across a broad range of realistic scenarios.

### v6

@@ Appendix B Projection algorithms
As discussed in the main text, the algorithm minimizes a penalized objective function along a decreasing sequence of scaling coefficients $\mu^{(1)}>\cdots>\mu^{(T)}$. At each stage, the solution $W^{(t)}$ serves as a warm start for the next problem at $\mu^{(t+1)}$. This continuation scheme produces a sequence of unconstrained matrices $\{W^{(t)}\}_{t=1}^{T}$ that converges as $\mu\to 0$ to a point satisfying the constraint $h(W)=0$. In our implementation, we set $\mu^{(1)}=1$, $\mu^{(t+1)}=\mu^{(t)}/2$, $T=10$, and $W^{(0)}=0\in\mathbb{W}$.
Algorithm 3 projects the acyclic matrix $\hat{W}$ from Algorithm 2 onto the $\ell_{1}$ ball of size $\lambda$.
@@
[…]
@@
The algorithm sorts the entries of $\hat{W}$ by absolute value, computes a threshold $\theta$ from the largest entries consistent with the $\ell_{1}$ constraint, and then applies soft-thresholding to shrink them toward zero (setting some exactly to zero) while preserving signs. The resulting matrix $W$ satisfies $\|W\|_{\ell_{1}}\leq\lambda$ by construction.
@@ Appendix E Implementation
  • DiBS and DiBS+: do not include a sparsity parameter;
@@
The experiments are run on a Linux workstation with an AMD Ryzen Threadripper PRO 5995WX CPU, 256GB RAM, and 2 × NVIDIA GeForce RTX 4090 GPUs. Each method is allocated either a single GPU or a single CPU core, with the experiments run in parallel across GPUs/CPU cores.
@@ Appendix F Simulation design
The synthetic datasets generated from linear DAGs are constructed as follows. We first randomly sample an Erdős–Rényi or scale-free graph with $p$ nodes and $s$ edges. The edges of this undirected graph are then oriented according to a randomly generated topological ordering. Each directed edge in the graph is then weighted by sampling weights uniformly on $0.7,-0.3]\cup[0.3,0.7]$, yielding a weighted adjacency matrix $W$. We then draw iid noise $\varepsilon_{1},\dots,\varepsilon_{n}$ and generate the variables as
@@
For synthetic datasets generated from nonlinear DAGs, we likewise sample an Erdős–Rényi or scale-free graph and orient it according to a random topological ordering. We then generate a feedforward neural network $f_{k}$ with a single hidden layer of 10 neurons for $k=1,\dots,p$. The weights of $f_{k}$ are sampled uniformly from $0.7,-0.3]\cup[0.3,0.7]$, with hidden layer weights of node $k$’s non-parents set to zero. The activation is a ReLU. The noise $\varepsilon_{1},\dots,\varepsilon_{n}$ is drawn iid and the variables are taken as
@@ Appendix G Additional experiments
G.1 Alternative simulation designs
We investigate the robustness of ProDAG to several alternative simulation designs. Figures 7 and 8 report results on dense graphs containing two and three times as many edges as nodes.
@@ G.1 Alternative simulation designs
Figures 9 and 10 report results on scale-free graphs whose degree exponent is set to two and three.
@@
Figures 11 and 12 report results from Gumbel noise distributions, while Figures 13 and 14 report results from exponential noise distributions.
Figure 12: Performance on synthetic datasets generated from linear Erdős–Rényi DAGs with $p=100$ nodes, $s=200$ edges, and Gumbel noise. The averages (solid points) and standard errors (error bars) are measured over 10 independently and identically generated datasets.
Figure 14: Performance on synthetic datasets generated from linear Erdős–Rényi DAGs with $p=100$ nodes, $s=200$ edges, and exponential noise. The averages (solid points) and standard errors (error bars) are measured over 10 independently and identically generated datasets.
Finally, Figure 15 report results from Gaussian noise distributions with heteroscedastic variance.
Figure 15: Performance on synthetic datasets generated from linear Erdős–Rényi DAGs with $p=20$ nodes, $s=40$ edges, and heteroscedastic Gaussian noise. The averages (solid points) and standard errors (error bars) are measured over 10 independently and identically generated datasets.
A Gaussian likelihood is used consistently across settings. Under all simulation designs, the relative performance of ProDAG is roughly similar to that in the main experiments, indicating that ProDAG remains reliable across a broad range of realistic scenarios.
G.2 Semi-synthetic datasets
We generate semi-synthetic datasets using the MUNIN (subnetwork #1) DAG, a large medical diagnostic network from the Bayesian Network Repository (https://www.bnlearn.com/bnrepository). The graph, which contains $p=186$ nodes and $s=273$ edges, is used to generate semi-synthetic datasets by sampling edge weights and noise in the same fashion as the fully synthetic experiments. Table 2 reports results with sample size $n=100$, showing that ProDAG performs strongly with the lowest Brier score and expected SHD.
Table 2: Performance on semi-synthetic datasets of size $n=100$ generated from the MUNIN graph of Sachs et al. (2005) with $p=186$ nodes, $s=273$ edges, and Gaussian noise. The averages and standard errors are measured over 10 independently and identically generated datasets. The best value of each metric is indicated in bold.
Brier score Exp. SHD Exp. F1 score (%) AUROC (%)
ProDAG $\mathbf{146\pm 4.0}$ $\mathbf{201\pm 4.9}$ $37\pm 2.0$ $87\pm 0.6$
DAGMA $305\pm 65.8$ $274\pm 64.9$ $\mathbf{67\pm 4.1}$ $\mathbf{89\pm 0.5}$
DiBS - - - -
DiBS+ - - - -
BayesDAG $279\pm 0.3$ $557\pm 0.9$ $1\pm 0.0$ $53\pm 0.7$
BOSS - - - -
While DAGMA scores well on expected F1 score and AUROC, ProDAG provides superior calibration and structure recovery. To our knowledge, this experiment represents one of the largest Bayesian DAG evaluations reported to date. Aside from ProDAG, the only methods to successfully run are DAGMA and BayesDAG. The other baselines encounter difficulties: DiBS and DiBS+ produce nonsensical results indicative of convergence issues, while BOSS crashes.
G.3 MCMC baseline
We compare ProDAG with Gadget (Viinikka et al., 2020), an earlier Bayesian approach that learns posteriors over DAGs via MCMC. We focus on linear Erdős–Rényi DAGs with $p=20$ nodes and $s=40$ edges, as Gadget does not scale to our larger setting. Table 3 reports results with sample size $n=100$, showing that Gadget substantially underperforms ProDAG.
Table 3: Performance on synthetic datasets of size $n=100$ generated from linear Erdős–Rényi DAGs with $p=20$ nodes, $s=40$ edges, and Gaussian noise. The averages and standard errors are measured over 10 independently and identically generated datasets. The best value of each metric is indicated in bold.
Brier score Exp. SHD Exp. F1 score (%) AUROC (%)
ProDAG $\mathbf{13\pm 1.8}$ $\mathbf{14\pm 1.5}$ $\mathbf{78\pm 2.7}$ $\mathbf{91\pm 1.6}$
Gadget $52\pm 1.8$ $37\pm 0.5$ $17\pm 1.5$ $65\pm 3.0$
This underperformance is not due to Gadget’s configuration, as we evaluated many settings and report the best-performing one.
G.4 Timings
We compare the run times of ProDAG with those of the Bayesian baselines in our main experiments. Table 4 reports timings on ($p=20$ nodes, $s=40$ edges) and nonlinear ($p=10$ nodes, $s=20$ edges) Erdős–Rényi DAGs with Gaussian noise.
Table 4: Run times in seconds with a sample size $n=100$ for linear ($p=20$ nodes) and nonlinear ($p=10$ nodes) settings. The averages and standard errors are measured over 10 independently and identically generated datasets. The best value in each setting is indicated in bold.
Linear Nonlinear
ProDAG $30\pm 7.0$ $61\pm 3.8$
DiBS $\mathbf{8\pm 0.7}$ $\mathbf{40\pm 0.4}$
BayesDAG $57\pm 1.1$ $122\pm 5.5$
The timings show that ProDAG is faster than BayesDAG and slower than DiBS. However, ProDAG’s superior performance in uncertainty quantification and structure recovery makes it a strong choice, even with the higher computational cost over DiBS.